# Learning parameter dependence for Fourier-based option pricing with tensor networks

### v3

@@ I Introduction
In light of this, applications of quantum-inspired algorithms, that is tensor network (TN) algorithms, to option pricing have also been studied as solutions in the present or the near future [5, 6, 7, 8]. Among them, this paper focuses on the application of tensor train (TT) [9] learning to the Fourier transform (FT)-based option pricing method [10, 11], following the original proposal in [6]. This option pricing method is based on converting the integration for the expected payoff in the space of the asset prices $\vec{S}(T)$ to that in the Fourier space, namely, the space of $\vec{z}$, the wavenumbers corresponding to the logarithm of $\vec{S}(T)$. After this conversion, the numerical integration is done more efficiently in many cases. Unfortunately, as is common in numerical integration of multivariate functions, this approach suffers from the curse of dimensionality: the FT-based method is efficient for single-asset options but for multi-asset options, its computational complexity increases exponentially. On the other hand, tensor network is the technique originally developed in quantum many-body physics to express state vectors with exponentially large dimension efficiently [^3] See [12, 13] as reviews. and recently, it has also been utilized in various fields such as machine learning [14, 15], quantum field theory [16, 17], and differential equations [7, 18]. Ref. [6] proposed to leverage the ability of tensor network to compress data as a high-dimensional tensor in order to express the functions of $\vec{z}$ involved in FT-based option pricing. The authors built TTs, a kind of tensor network, approximating those functions by a TT learning algorithm called tensor cross interpolation (TCI) [19, 20, 17, 21] and evaluated the integral involving them efficiently, which led to the significant speed-up of FT-based option pricing in their test cases.
@@
To address these issues, we focus on another usage of tensor networks that can embed parameter dependence [22, 23, 24] to make FT-based option pricing more efficient. Namely, we learn TTs that approximate the functions including not only the dependence on $\vec{z}$ but also that on parameters in the asset price model such as the volatilities and the present asset prices, by a single application of TCI for each function. We use these tensor networks to perform fast option pricing in response to various parameter changes. To evaluate this approach, we consider two scenarios as benchmarks, varying volatilities and present stock prices. In the test cases, it is seen that for up to 11 assets, the computational complexity of our proposed method, which is measured by the number of elementary operations is advantageous to that of the Monte Carlo method with $10^{5}$ paths, a common setting in practice. Indeed, the computational complexity of our proposed method is better than that of Monte Carlo by a factor of $O(10^{4})$ for $d<10$ and by a factor of $O(10^{3})$ for about $d=10$ or $11$ (refer to FIG. 2). We also confirm numerically that the accuracy of our method is within the statistical error in the Monte Carlo method with $10^{5}$ paths. In summary, these results suggest that at least in some cases, our proposed method has advantages over Monte Carlo in terms of computational complexity, keeping the accuracy.
@@ II Tensor train
|$\ket{F_{x_{1},x_{2},\ldots,x_{d}}}$
$=\sum_{l_{1}}^{\chi_{1}}\cdots\sum_{l_{d-1}}^{\chi_{d}}F_{l_{1},x_{1}}^{(1)}F_{l_{1}l_{2},x_{2}}^{(2)}\cdots F_{l_{d-1},x_{d}}^{(d)}\ket{\vec{x}},$ (2)
@@ II.1 Compression tenchniques
II.1.1 Tensor cross interpolation
@@
$\epsilon_{\mathrm{max}}=\frac{\|F_{\mathrm{TT}}-\tilde{F}_{\mathrm{TT}}\|_{\mathrm{max}}}{\|F_{\mathrm{TT}}\|_{\mathrm{max}}}$ (3)
where $F_{\mathrm{TT}}$ is a target tensor, $\tilde{F}_{\mathrm{TT}}$ is a low-rank approximation, and the maximum norm is evaluated as the maximum of the absolute values of the entries at the pivots selected already. The computational complexity of TCI is roughly proportional to the number of elements in the TT, which is $O(d\chi^{2}N)$ with $\chi_{1},\cdots,\chi_{d}$ fixed to $\chi$. In addition, considering the case that zero is included in the reference function value, the error should be normalized by $\|F_{\mathrm{TT}}\|_{\mathrm{max}}$.
II.1.2 Singular value decomposition
In this study, we use singular value decomposition (SVD) to compress further the TTs obtained by TCI with its error threshold $\epsilon_{\mathrm{TCI}}$ set to a sufficiently low. The original TT, denoted as $F_{\mathrm{TT}}$, can be approximated by another TT, $\tilde{F}_{\mathrm{TT}}$, with reduced (optimized) bond dimensions such that
@@ II.1.2 Singular value decomposition
III Fourier transform-based option pricing aided by tensor cross interpolation
III.1 Fourier transform-based option pricing
@@
$V(\vec{p})$
$=\frac{e^{-rT}}{2\pi}\sum_{j_{1},\cdots,j_{d}=-N/2}^{N/2}\phi(-\vec{z}_{\mathrm{gr},\vec{j}}-i\vec{\alpha})\hat{v}(\vec{z}_{\mathrm{gr},\vec{j}}+i\vec{\alpha})\Delta_{\rm vol}.$ (13)
@@
where $\eta_{m}$ is the integration step size in the $m$-th direction. $\eta_{m}$ is a hyperparameter that must be appropriately determined according to the parameter $\vec{p}$ and thus written as $\eta_{m}(\vec{p})$ in (14). In fact, in our demonstration, it has been numerically observed that the optimal value of $\eta$ changes abruptly in response to variations in $\sigma$ within a small range; see Sec. V for the detail. $\Delta_{\rm vol}:=\Pi_{m=1}^{d}\eta_{m}$ is the volume element.
III.2 Fourier transform-based option pricing with tensor trains
@@
III.3 Monte Carlo-based option pricing
@@
IV Learning parameter dependence with tensor networks
IV.1 Outline
@@
$K^{k_{1},\ldots,k_{d}}_{k_{1},\ldots,k_{d}}=\sum^{N/2}_{j_{1},\ldots,j_{d}=-N/2}\tilde{\phi}^{k_{1},\ldots,k_{d}}_{j_{1},\ldots,j_{d}}\tilde{v}^{j_{1},\ldots,j_{d}}_{k_{1},\ldots,k_{d}},$ (17)
and optimize the bond dimensions of this resultant MPO $K^{k_{1},\ldots,k_{d}}_{k_{1},\ldots,k_{d}}$ using SVD. Having this MPO, we get the option price for the specified value of $\vec{p}$ as illustrated in Fig. 1 (b). By fixing the index $\vec{k}$ of $K^{k_{1},\ldots,k_{d}}_{k_{1},\ldots,k_{d}}$ as $\hat{k}_{1},\ldots,\hat{k}_{d}$ to the value corresponding to the specified $\vec{p}$, we get the option price for the specified $\vec{p}$ as
$V(\vec{p}_{\mathrm{gr},\vec{\hat{k}}})\simeq\frac{e^{-rT}}{2\pi}K^{\hat{k}_{1},\ldots,\hat{k}_{d}}_{\hat{k}_{1},\ldots,\hat{k}_{d}}\Delta_{\mathrm{vol}}.$ (18)
@@
We also mention the ordering of the local indices of the TT. Two core tensors in the TT that correspond to $z_{j}$ and $p_{j}$ of the same asset are arranged next to each other, i.e., the order is ($z_{1}p_{1}z_{2}p_{2}\cdots z_{d}p_{d}$). In the demonstration in Sec. V, we have numerically found that this arrangement allows us to compress the TTs with parameter dependence while maintaining the accuracy of the option pricing. On the other hand, if the core tensors on $\vec{z}$ and those on $\vec{p}$ are completely separated, i.e.,($z_{1}z_{2}\cdots z_{d}p_{1}p_{2}\cdots p_{d}$), we have found that the accuracy of the option pricing get worse since TCI fails to learn this tensor trains. However, the optimal arrangement of the local indices may vary, for example, depending on the correlation matrix: intuitively, the core tensors corresponding to highly correlated assets should be placed nearby. Although we do not discuss it in detail, this is an important topic for future research.
In the two test cases for our proposed method in Sec. V, we will identify $\vec{p}$ as the volatility $\vec{\sigma}$ or the present asset price $\vec{S}_{0}$, the varying market parameters that particularly affect the option prices. Note that it is possible to include dependencies on other parameters in the TTs. For example, taking into account the dependence on the parameters concerning the option contract, such as the maturity $T$ and the strike $K$, enables us to price different option contracts with a single set of TTs. Although this is a promising approach, there might be some issues. For example, it is non-trivial whether TTs incorporating many parameter dependencies have a low-rank structure. Thus, we will leave such a study for future work.
@@
IV.2 Computational complexity
The computational complexity of TT-based option pricing, involving computation for obtaining the specific tensor components $K^{\hat{k}_{1},\ldots,\hat{k}_{d}}_{\hat{k}_{1},\ldots,\hat{k}_{d}}$ for the fixed values of $\hat{k}_{1},\ldots,\hat{k}_{d}$ is $O(d\chi_{K}^{2})$ [25]. Here, we denote the maximum bond dimensions as $\chi_{K}$ In fact, the bond dimension depends on the bond index, and it is necessary to account for this for an accurate evaluation of the number of operations. Indeed, we consider this point in evaluating the computational complexity of TT-based option pricing demonstrated in Sec. V.
Here, we ignore the computational complexity of all the processes in FIG. 1 (a) and consider that in FIG. 1 (b) after we get the MPO only. This is reasonable if we can use plenty of time to perform these tasks before we need fast option pricing. As discussed in Sec. VI, we can reasonably find such a situation in practice.
V Numerical demonstration
V.1 Details
@@
The TCI code has not yet been made publicly available. Marc K. Ritter, Hiroshi Shinaoka, and Jan von Delft provided us with access to their TCI code. ITensors.jl [29] was utilized for computing the contraction and inner product of the tensor trains. The Monte Carlo simulations were carried out using tf-quant-finance [30]. Parallelization was not employed in either case. GPUs were not utilized in any of the calculations. The computations were performed on a 2019 MacBook Pro featuring a 2.3GHz 8-core Intel Core i9 processor and 16GB of 2667MHz DDR4 memory.
V.2 Results
@@
The results are summarized in TABLE 2. In particular, the computational complexity versus $d$ is plotted in FIG. 2. The maximum relative error of TT-based option pricing among runs for 100 random parameter sets is represented by $r_{\mathrm{TT}}$. The average relative error and standard deviation of the 20 runs of the MC-based method for the same parameter sets are denoted by $r_{\mathrm{MC}}$ and $\sigma_{\text{MC}}$, respectively. The computational complexities of TT-based and MC-based option pricing are represented by $c_{\mathrm{TT}}$ and $c_{\mathrm{MC}}$, respectively. Also, the computational time of TT-based and MC-based option pricing are denoted as $t_{\mathrm{TT}}$ and $t_{\mathrm{MC}}$, respectively. To maintain the desired accuracy of option pricing, we set the tolerance of TCI sufficiently low, concretely $\epsilon_{\mathrm{TCI}}=10^{-9}$ for $d\leq 10$, and $\epsilon_{\mathrm{TCI}}=10^{-10}$ for $d=11$. Subsequently, we reduce the bond dimension by SVD, with the tolerance of SVD set to $\epsilon_{\mathrm{SVD}}^{\phi}$ and $\epsilon_{\mathrm{SVD}}^{\hat{v}}$ in TABLE 2 for $\phi$ and $\hat{v}$, respectively. The maximum bond dimensions of the MPOs for $\phi$ and $\hat{v}$ are denoted as $\chi_{\phi}$ and $\chi_{\hat{v}}$, respectively, in TABLE 2. We set the tolerance of SVD for obtaining $\chi_{K}$, maximum bond dimension of $K^{k_{1},\ldots,k_{d}}_{k_{1},\ldots,k_{d}}$, set to $\epsilon^{K}_{\mathrm{SVD}}=10^{-10}$ regardless of the number of assets to maintain the accuracy.
V.2.1 The case of varying $\vec{\sigma}$
TABLE 2 (a) shows the computational results of TT-based option pricing when we consider $\vec{\sigma}$ dependence. TT-based option pricing demonstrates advantages in terms of computational complexity and time over the MC-based method. The bond dimensions of the TT results for $d=5$ and $d=10$ are depicted in Fig. 3(a). By applying SVD to the TTs trained via TCI, the bond dimensions between tensors related to $p_{m}$ and $z_{m+1}$ could generally be maintained at less than 10. The details of this compression by SVD are described in Appendix A. The maximum bond dimension $\chi_{K}$ of $K^{k_{1},\ldots,k_{d}}_{k_{1},\ldots,k_{d}}$ is maintained at a small value of about 5, when $\eta$ is constant, but increases to about 20 when $\eta$ is automatically determined for each asset, which leads to increased computational complexity at around $d=10$. Still, the computational complexity is much lower than that of MC-based option pricing. The accuracy of the TT-based option pricing fell within the error range of the Monte Carlo method with $10^{5}$ paths.

### v4

@@ 1 Introduction
In light of this, applications of quantum-inspired algorithms, that is tensor network (TN) algorithms, to option pricing have also been studied as solutions in the present or the near future [5, 6, 7, 8]. Among them, this paper focuses on the application of tensor train (TT) [9] learning to the Fourier transform (FT)-based option pricing method [10, 11], following the original proposal in [6]. This option pricing method is based on converting the integration for the expected payoff in the space of the asset prices $\vec{S}(T)$ to that in the Fourier space, namely, the space of $\vec{z}$, the wavenumbers corresponding to the logarithm of $\vec{S}(T)$. After this conversion, the numerical integration is done more efficiently in many cases. Unfortunately, as is common in numerical integration of multivariate functions, this approach suffers from the curse of dimensionality: the FT-based method is efficient for single-asset options but for multi-asset options, its computational complexity increases exponentially. On the other hand, tensor network is the technique originally developed in quantum many-body physics to express state vectors with exponentially large dimension efficiently [^3] See [12, 13] as reviews. and recently, it has also been utilized in various fields such as machine learning [14, 15], quantum field theory [16, 17, 18], and partial differential equations [7, 19]. Ref. [6] proposed to leverage the ability of tensor network to compress data as a high-dimensional tensor in order to express the functions of $\vec{z}$ involved in FT-based option pricing. The authors built TTs, a kind of tensor network, approximating those functions by a TT learning algorithm called tensor cross interpolation (TCI) [20, 21, 17, 22] and evaluated the integral involving them efficiently, which led to the significant speed-up of FT-based option pricing in their test cases.
@@
To address this issue, we focus on another usage of tensor networks that can embed parameter dependence [23, 24, 25] to make FT-based option pricing more efficient. Namely, we learn TTs that approximate the functions including not only the dependence on $\vec{z}$ but also that on parameters in the asset price model such as the volatilities and the present asset prices, by a single application of TCI for each function. We use these tensor networks to perform fast option pricing in response to various parameter changes. To evaluate this approach, we consider two scenarios as benchmarks, varying volatilities and present stock prices. In the test cases, it is seen that for up to 11 assets, the computational complexity of our proposed method, which is measured by the number of elementary operations is advantageous to that of the Monte Carlo method with $10^{5}$ paths, a common setting in practice. Indeed, the computational complexity of our proposed method is better than that of Monte Carlo by a factor of $O(10^{4})$ for $d<10$ and by a factor of $O(10^{3})$ for about $d=10$ or $11$ (refer to FIG. 2). We also confirm numerically that the accuracy of our method is within the statistical error in the Monte Carlo method with $10^{5}$ paths. In summary, these results suggest that at least in some cases, our proposed method has advantages over Monte Carlo in terms of computational complexity, keeping the accuracy.
@@ 2 Tensor train
|$\ket{F_{x_{1},x_{2},\ldots,x_{d}}}=\sum_{l_{1}}^{\chi_{1}}\cdots\sum_{l_{d-1}}^{\chi_{d}}F_{l_{1},x_{1}}^{(1)}F_{l_{1}l_{2},x_{2}}^{(2)}\cdots F_{l_{d-1},x_{d}}^{(d)}\ket{\vec{x}},$ (2)
@@ 2.1 Compression tenchniques
2.1.1 Tensor cross interpolation
@@
$\epsilon_{\mathrm{max}}=\frac{\|F_{x_{1},x_{2},\ldots,x_{d}}-\tilde{F}_{\mathrm{TT}}\|_{\mathrm{max}}}{\|F_{x_{1},x_{2},\ldots,x_{d}}\|_{\mathrm{max}}}$ (3)
where $F_{x_{1},x_{2},\ldots,x_{d}}$ is a target tensor [^5] Here, we assume that we can access the arbitrary elements of the target tensor. Note that we do not need to store all the elements of the target tensor., $\tilde{F}_{\mathrm{TT}}$ is a low-rank approximation, and the maximum norm is evaluated as the maximum of the absolute values of the entries at the pivots selected already. The computational complexity of TCI is roughly proportional to the number of elements in the TT, which is $O(d\chi^{2}N)$ with $\chi_{1},\cdots,\chi_{d}$ fixed to $\chi$. In addition, considering the case that zero is included in the reference function value, the error should be normalized by $\|F_{x_{1},x_{2},\ldots,x_{d}}\|_{\mathrm{max}}$.
2.1.2 Singular value decomposition
In this study, we use singular value decomposition (SVD) to compress further the TTs obtained by TCI with its error threshold $\epsilon_{\mathrm{TCI}}$ set to a sufficiently low. The TT obtained from TCI, denoted as $F_{\mathrm{TT}}$, can be approximated by another TT, $\tilde{F}_{\mathrm{TT}}$, with reduced (optimized) bond dimensions such that
@@ 2.1.2 Singular value decomposition
3 Fourier transform-based option pricing aided by tensor cross interpolation
3.1 Fourier transform-based option pricing
@@
$V(\vec{p})=\frac{e^{-rT}}{2\pi}\sum_{j_{1},\cdots,j_{d}=-N/2}^{N/2}\phi(-\vec{z}_{\mathrm{gr},\vec{j}}-i\vec{\alpha})\hat{v}(\vec{z}_{\mathrm{gr},\vec{j}}+i\vec{\alpha})\Delta_{\rm vol}.$ (13)
@@
where $\eta_{m}$ is the integration step size in the $m$-th direction. $\eta_{m}$ is a hyperparameter that must be appropriately determined according to the parameter $\vec{p}$ and thus written as $\eta_{m}(\vec{p})$ in (14). In fact, in our demonstration, it has been numerically observed that the optimal value of $\eta$ changes abruptly in response to variations in $\sigma$ within a small range; see Sec. 5 for the detail. $\Delta_{\rm vol}:=\Pi_{m=1}^{d}\eta_{m}$ is the volume element.
3.2 Fourier transform-based option pricing with tensor trains
@@
3.3 Monte Carlo-based option pricing
@@
4 Learning parameter dependence with tensor networks
4.1 Outline
@@
$O^{k_{1},\ldots,k_{d}}_{k_{1},\ldots,k_{d}}=\sum^{N/2}_{j_{1},\ldots,j_{d}=-N/2}\tilde{\phi}^{k_{1},\ldots,k_{d}}_{j_{1},\ldots,j_{d}}\tilde{v}^{j_{1},\ldots,j_{d}}_{k_{1},\ldots,k_{d}},$ (17)
and optimize the bond dimensions of this resultant MPO $O^{k_{1},\ldots,k_{d}}_{k_{1},\ldots,k_{d}}$ using SVD. Having this MPO, we get the option price for the specified value of $\vec{p}$ as illustrated in Fig. 1 (b). By fixing the index $\vec{k}$ of $O^{k_{1},\ldots,k_{d}}_{k_{1},\ldots,k_{d}}$ as $\hat{k}_{1},\ldots,\hat{k}_{d}$ to the value corresponding to the specified $\vec{p}$, we get the option price for the specified $\vec{p}$ as
$V(\vec{p}_{\mathrm{gr},\vec{\hat{k}}})\simeq\frac{e^{-rT}}{2\pi}O^{\hat{k}_{1},\ldots,\hat{k}_{d}}_{\hat{k}_{1},\ldots,\hat{k}_{d}}\Delta_{\mathrm{vol}}.$ (18)
@@
We also mention the ordering of the local indices of the TT. Two core tensors in the TT that correspond to $z_{j}$ and $p_{j}$ of the same asset are arranged next to each other, i.e., the order is ($z_{1}p_{1}z_{2}p_{2}\cdots z_{d}p_{d}$). In the demonstration in Sec. 5, we have numerically found that this arrangement allows us to compress the TTs with parameter dependence while maintaining the accuracy of the option pricing. On the other hand, if the core tensors on $\vec{z}$ and those on $\vec{p}$ are completely separated, i.e.,($z_{1}z_{2}\cdots z_{d}p_{1}p_{2}\cdots p_{d}$), we have found that the accuracy of the option pricing get worse since TCI fails to learn this tensor trains. However, the optimal arrangement of the local indices may vary, for example, depending on the correlation matrix: intuitively, the core tensors corresponding to highly correlated assets should be placed nearby. Although we do not discuss it in detail, this is an important topic for future research.
In the two test cases for our proposed method in Sec. 5, we will identify $\vec{p}$ as the volatility $\vec{\sigma}$ or the present asset price $\vec{S}_{0}$, the varying market parameters that particularly affect the option prices. Note that it is possible to include dependencies on other parameters in the TTs. For example, taking into account the dependence on the parameters concerning the option contract, such as the maturity $T$ and the strike $K$, enables us to price different option contracts with a single set of TTs. Although this is a promising approach, there might be some issues. For example, it is non-trivial whether TTs incorporating many parameter dependencies have a low-rank structure. Thus, we will leave such a study for future work.
@@
4.2 Computational complexity
The computational complexity of TT-based option pricing, involving computation for obtaining the specific tensor components $O^{\hat{k}_{1},\ldots,\hat{k}_{d}}_{\hat{k}_{1},\ldots,\hat{k}_{d}}$ for the fixed values of $\hat{k}_{1},\ldots,\hat{k}_{d}$ is $O(d\chi_{O}^{2})$ [26]. Here, we denote the maximum bond dimensions as $\chi_{K}$ In fact, the bond dimension depends on the bond index, and it is necessary to account for this for an accurate evaluation of the number of operations. Indeed, we consider this point in evaluating the computational complexity of TT-based option pricing demonstrated in Sec. V.
Here, we ignore the computational complexity of all the processes in FIG. 1 (a) and consider that in FIG. 1 (b) after we get the MPO only. This is reasonable if we can use plenty of time to perform these tasks before we need fast option pricing. As discussed in Sec. 6, we can reasonably find such a situation in practice.
5 Numerical demonstration
5.1 Details
@@
TensorCrossInterpolation.jl for TCI [30] has not yet been made publicly available. Marc K. Ritter, Hiroshi Shinaoka, and Jan von Delft provided us with access to this. ITensors.jl [31] was utilized for computing the contraction and inner product of the tensor trains. The Monte Carlo simulations were carried out using tf-quant-finance [32]. Parallelization was not employed in either case. GPUs were not utilized in any of the calculations. The computations were performed on a 2019 MacBook Pro featuring a 2.3GHz 8-core Intel Core i9 processor and 16GB of 2667MHz DDR4 memory.
5.2 Results
@@
The results are summarized in TABLE 2. In particular, the computational complexity versus $d$ is plotted in FIG. 2. The maximum relative error of TT-based option pricing among runs for 100 random parameter sets is represented by $r_{\mathrm{TT}}$. The average relative error and standard deviation of the 20 runs of the MC-based method for the same parameter sets are denoted by $r_{\mathrm{MC}}$ and $\sigma_{\text{MC}}$, respectively. The computational complexities of TT-based and MC-based option pricing are represented by $c_{\mathrm{TT}}$ and $c_{\mathrm{MC}}$, respectively. Also, the computational time of TT-based and MC-based option pricing are denoted as $t_{\mathrm{TT}}$ and $t_{\mathrm{MC}}$, respectively. To maintain the desired accuracy of option pricing, we set the tolerance of TCI sufficiently low, concretely $\epsilon_{\mathrm{TCI}}=10^{-9}$ for $d\leq 10$, and $\epsilon_{\mathrm{TCI}}=10^{-10}$ for $d=11$. Subsequently, we reduce the bond dimension by SVD, with the tolerance of SVD set to $\epsilon_{\mathrm{SVD}}$ in TABLE 2 for $\phi$ and $\hat{v}$, respectively. The maximum bond dimensions of the MPOs for $\phi$ and $\hat{v}$ are denoted as $\chi_{\phi}$ and $\chi_{\hat{v}}$, respectively, in TABLE 2. We set the tolerance of SVD for obtaining $\chi_{K}$, maximum bond dimension of $O^{k_{1},\ldots,k_{d}}_{k_{1},\ldots,k_{d}}$, set to $\epsilon^{O}_{\mathrm{SVD}}=10^{-10}$ regardless of the number of assets to maintain the accuracy.
5.2.1 The case of varying $\vec{\sigma}$
TABLE 2 (a) shows the computational results of TT-based option pricing when we consider $\vec{\sigma}$ dependence. TT-based option pricing demonstrates advantages in terms of computational complexity and time over the MC-based method. The bond dimensions of the TT results for $d=5$ and $d=10$ are depicted in Fig. 3(a). By applying SVD to the TTs trained via TCI, the bond dimensions between tensors related to $p_{m}$ and $z_{m+1}$ could generally be maintained at less than 10. The details of this compression by SVD are described in Appendix A. The maximum bond dimension $\chi_{O}$ of $O^{k_{1},\ldots,k_{d}}_{k_{1},\ldots,k_{d}}$ is maintained at a small value of about 5, when $\eta$ is constant, but increases to about 20 when $\eta$ is automatically determined for each asset, which leads to increased computational complexity at around $d=10$. Still, the computational complexity is much lower than that of MC-based option pricing. The accuracy of the TT-based option pricing fell within the error range of the Monte Carlo method with $10^{5}$ paths.